Mendelian Genetics problems:

1)

What are the probabilities of a particular genotype or phenotype for an offspring for the following parental genotypes?

a.     Monohybrid cross if one parent is homozygous recessive and one parent is homozygous recessive? What if both are heterozygotes?

b.     Dihybrid cross if one parent is homozygous recessive and one parent is homozygous recessive? What if both are heterozygotes?

2)

What are the probabilities that two parents with the genotypes AaBbccDD x AabbCcDd will have an offspring with the genotype AAbbCcDD?

What about the phenotype aBcD?
Answer #1

Answer:

1a).

If both the parents are homozygous recessive, all progeny are homozygous.

If both the parents are heterozygous,

Dominant phenotype (3) : Recessive phenotype (1)

Homozygous dominant genotype (1) : Heterozygous dominant genotype (2): Homozygous recessive genotype (1)

1b).

If both the parents are homozygous recessive, all progeny are homozygous.

aabb x aabb----------parents

aabb----offspring

If both the parents are heterozygous, AaBb x AaBb---parents

A_B_ = 9

A_bb = 3

aa B_ = 3

aabb=1

2) AaBbccDD x AabbCcDd

Aa x Aa = AA(1/4), Aa(1/2) & aa(1/4)

Bb x bb = Bb (1/2) & bb(1/2)

cc x Cc = Cc (1/2) & cc (1.2)

DD x Dd = DD (1/2) & Dd (1/2)

The probability of AAbbCcDD genotype= 1/4 x 1/2 x 1/2 x 1/2 =1/32

The phenotype aBcD = aaB_ccD_ = 1/4 x 1/2 x 1/2 x 1= 1/16